OpenStudy (anonymous):

A spherical balloon is filled with helium at the rate of 8 in^3/sex . How fast is the diameter of the balloon increasing when the volume of the ballon is 36p in ^3?

OpenStudy (anonymous):

id assume 4

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

relationship of 36 and 8

OpenStudy (anonymous):

between*

OpenStudy (anonymous):

4.5

OpenStudy (surry99):

you need to apply the chain rule of calculus

OpenStudy (anonymous):

how

OpenStudy (surry99):

What is the expression for the volume of a sphere?

OpenStudy (anonymous):

4/3 pie r

OpenStudy (anonymous):

r 3d

OpenStudy (anonymous):

r to the 3

OpenStudy (anonymous):

4/3 pi r3

OpenStudy (surry99):

good, now express it in terms of the diameter instead of the radius

OpenStudy (anonymous):

I am so confused

OpenStudy (anonymous):

so 4/3 pi 2r3

OpenStudy (anonymous):

because the diam is twice as large as radius

OpenStudy (anonymous):

so the diameter is 4/3 pi 2r3

OpenStudy (anonymous):

yea

OpenStudy (surry99):

V = 4/3*pi*r^3 but r=d/2 Now substitute and simplify

OpenStudy (anonymous):

ok but how fast is the diameter of the balloon increasing when the volume of the ballon is 36pi

OpenStudy (anonymous):

2/3

OpenStudy (surry99):

Can you solve for V in terms of d?